# Operations & Supply Chain Management: Principles and Practice

### v1

@@ 3 Supply Chain Management
Similarly, the rapid growth of e-commerce has introduced distinct operational complexities, such as the need for efficient last-mile delivery and inventory management in a digital marketplace (Janjevic and Winkenbach, 2020). In the context of e-commerce, Calzavara et al. (2023) investigated different strategies for e-grocery logistics, comparing the costs and performances of various network designs to determine the most efficient approach in response to fluctuating online demand. Another paper (Agac et al., 2023) introduced a novel mathematical model using Mixed-Integer Nonlinear Programming for optimising blood supply chain networks, focusing on location, allocation, and routing. A variant of the Genetic Algorithm is applied, and the model’s effectiveness is demonstrated through a case study in the Eastern Black Sea Region of the Turkish Red Crescent.
@@
Purchasing and supply management can contribute to organisational efficiency, risk management and innovation (Axelsson and Håkansson, 1984). In financial terms, this translates into cost reduction, assets optimisation and revenue growth. Growing emphasis is placed on the contribution to non-financial objectives such as environmental and social sustainability and governance (Quarshie et al., 2016). The exact definition of purchasing and supply management – or, in short: procurement – has seen a shift in emphasis since the 1960s, when the field became subject of academic education and research (Wynstra et al., 2019). The first textbooks focused on differences with consumer buying behaviour and on the tactical and operational purchasing process (e.g., Webster and Wind, 1972). By the 1990s, definitions started to explicate the variety of buying needs, such as the one by Van Weele (1994) that distinguished between direct (Bill of Material) and indirect spend. In the 2000s, definitions put the spotlight on the identification, development and maintenance of relations with suppliers (e.g., Axelsson et al., 2005). Consequently, purchasing and supply management is now seen to govern both tactical and operational, more short-term and internally oriented decisions and actions (purchasing) and strategic, more long term and externally oriented decisions and actions (supply), as depicted in figure 4.
@@
The upper, strategic/tactical part of this process model is commonly referred to as ‘sourcing, and is the focus of this section. The bottom part of the model is referred to as ‘purchase to pay’.
@@
Specification of need pertains to both quantities and qualities. Especially for direct purchase items, need quantities are primarily derived from final demand forecasts (see §4.6). Specification of needs in terms of qualities is typically a cross-functional process, where the procurement department gathers information and preferences from internal customers and technical specialists. Also, current and potential suppliers can provide valuable inputs. Giving suppliers influence over the specifications of their offerings positively affects both the efficiency and effectiveness of new product development projects (Suurmond et al., 2020).
@@
In a growing number of organisations, procurement will have the possibility if not the imperative to step up from just purchasing goods and services to becoming a function of real end-to-end value entrepreneurs. Together with the growing digital technology support also for strategic processes, this will have a major impact on how procurement adds value to the organisation.
@@
The origins of academic study on the management of buyer-supplier relationships can be briefly traced to the history of economics and specifically contracting through the theories of agency, finance, property rights and stakeholder theory Jensen and Meckling (1976). Following the industrial revolution as firms moved towards ‘new’ mass production techniques in the late 1920s and early 1930s scholars began to study the idea of ‘what is a firm’ and think about where a firm begins and ends.
@@
This work was synthesised by the Industrial Marketing and Purchasing (IMP) Group in the early 1980s and 1990s under the leadership of Håkansson (1982) to develop an empirical analysis of firm networks focusing on the relationship interdependencies and typologies of interactions ranging from short-term episodic to long-term ‘deep’ and embedded interactions (Johnsen, 2018; Ellram and Ueltschy Murfield, 2019). Collaborative business relationships are viewed as value adding process within the network of organisations or actors. Building on this foundational work we suggest a working definition of buyer-supplier relationships as follows:
@@
The strategic importance of buyer-supplier relationships has been increasingly studied by scholars since the late 1970s and 1980s with the development of Kraljic’s (1983) seminal portfolio analysis tool, positioning spending based on its relative market availability (essentially a measure of risk exposure) and cost or value to suggest management strategies based on a four-quadrant model. Moreover, development in buyer-supplier relationship thinking closely correlates to changes in the economic business environment and particularly in the paradigm changes in manufacturing which can be described as moving from craft, mass, lean and agile (Lamming, 1993) and laterally towards mass-customisation. Combining the portfolio analysis of Kraljic and his analysis of sourcing risk with the literature on contracts, firm boundaries and competitive advantage, led scholars to observe and develop a variety of buyer-supplier positioning and relationship measure concepts, tools and techniques (Cousins and Spekman, 2003).
During this time the business environment moved from local/regional sourcing to international and then global sourcing. The success of Japan following World War II as a major manufacturer employing innovative and novel manufacturing techniques began to change the way Western manufacturers built, designed and sourced their products – moving from one-stop manufacturers to systems integrators, particularly true of the automotive and aerospace industries (Hines et al., 1999). Techniques such as Just-In-Time (JIT) and Total Quality Management (TQM) involved working and coordinating closely with suppliers (Cousins et al., 2006). This meant manufacturers had to think much more strategically about how they would work with suppliers to compete. They restructured their supply bases moving to Tier levels, creating fewer but higher value suppliers. Harland (1996)and Harland et al. (1999) discuss the move from supplier dyads, chains to networks which typifies this transition in relationship discourse. Moving from multiple to more sole sourced approaches inevitably moved them into the ‘bottleneck’ and ‘critical’ aspects of Kraljic’s portfolio, meaning that the level of risk increased but so did the potential value. The question then moved onto how can buyer-organisations maximise that value in an environment of what became known as ‘co-opetition’.
More recently, the rise of China (and since other parts of East Asia, India and currently Africa) as a low-cost manufacturing environment encouraged large manufacturing organisations to outsource manufacturing and source product globally. This allowed for low-cost production, but naturally this strategy comes with risks, which are known as supply chain disruptions and are exacerbated in supply chains exposed to critical geographic capability concentration, such as those found in the advanced electronics or semiconductor industries found in Taiwan and South Korea. As the complexity of the supply network increases so does the risk of disruption to supply.
@@
Over the last fifty or so years there has been a substantial discourse on all aspects of buyer-supplier relationships creating a rich and varied body of knowledge. It reasons that most problems and questions focused on buyer-supplier relationships are multi-faceted. For scholars in this area, before embarking on research it is important to consider the unit of analysis and then select an appropriate theoretical lens(es). Essentially the research in this area can be classified into three broad but inter-related theoretical areas, economic, strategic and behavioural. Firstly, the economic lens uses the theory of Coase (1937) and specifically Williamson (1985) which is TCE. The unit of analysis here is the transaction at product or service level and focuses primarily on the risks of manufacturing within the boundaries of the firm or outside of it – known as market vs hierarchy. The risks are known as transaction costs and contain a variety of characteristics that can potentially be modelled. That said, transaction costs are notoriously difficult to calculate. The second major theoretical lens developed by Prahalad and Hamel (2009) and extended by Barney (1991) is the Resource Based View (RBV) of the firm – the focus here is on competitive advantage at the level of the firm, creating and managing value added relationships becomes a strategic competitive advantage to the focal firm. The seminal work of Dyer and Singh (1998) in this area builds on the previous ideas of the IMP of networks and dyads and theorises the concept of the ‘relational’ view. This paper has been extended (Dyer et al., 2018) to consider the value creation and capture benefits of building alliances and collaborations, viewing this as a dynamic capability. The third and final theoretical lens is ‘social capital’. Social capital can broadly be defined as the network of relationships among people who live and work in a particular society or organisation. Tools such as Social Network Theory (SNA) facilitate the mapping of these network relationships and to understand the strength of these ties and how information and knowledge flows between them (Granovetter, 1983). Leveraging these networks within buyer-supplier relationships can add significant value to the buyer and supplier organisations by facilitating innovations, learning, new product development and cost reduction and process improvement programmes (Lawson et al., 2009). These three theoretical lenses are not mutually exclusive however scholars tend to select one meta theory, this can be further supported by a secondary theory approach aimed at explaining some of the more nuanced issues within the research question (McIvor, 2009).
@@
The current political, economic, social and technological business environment is extremely difficult to navigate, fraught with supply chain disruptions, aggressive competition, sustainability pressures and move towards de-globalisation and protectionism. It is further compounded with profound changes in the way new technologies such as Artificial Intelligence (AI), Blockchain and the analysis of Big Data are impacting business and supply chain operations. These technologies are creating greater visibility and traceability allowing organisations to analyse and adjust the way they do business offering positive as well as negative consequences (Gligor et al., 2021). Facilitating relationships through the analysis of big data sets, help to analyse and resolve quality problems and maintain standards and improve innovation and problem solving (Rahman et al., 2023). However, it can also drastically reduce personal interactions leading to a lack of personal trust, socialisation and knowledge exchange (Cousins and Menguc, 2006; Lawson et al., 2009). These relational bonds are essential for efficient and effective problem solving and innovation development and knowledge transfer (Lawson et al., 2008). The impact of new technology while offering many advantages it can also negatively impact interfirm relationships.
@@
Stakeholder management is a critical aspect of operations and supply chain management (OCSM). Contemporary economic activity is becoming increasingly complex and requires coordination and collaboration within vast networks of globally distributed actors involved in delivering products or services to the market (Co and Barro, 2009; Silvestre et al., 2023). The efficiency of the supply chain depends on successful coordination among diverse actors, such as customers, retailers, logistics providers, manufacturers, and governments, through processes such as forecasting, planning, inventory management, and delivery (Najjar and Yasin, 2023; Silvestre et al., 2023). In today’s relational global economy, organisational success depends on active relationship management. Recent geopolitical and pandemic-related disruptions have underscored the importance of these interdependencies (Devi et al., 2023; Montoya-Torres et al., 2023).
@@
Stakeholder management research emphasises the need to consider diverse interests in strategic decision making (Donaldson and Preston, 1995b; Freeman, 1984; Laplume et al., 2008; Parmar et al., 2010). Fundamental to stakeholder thinking is the suggestion that “if we adopt as a unit of analysis the relationships between a business and the groups and individuals who can affect or are affected by it, then we have a better chance to” create and distribute value to multiple actors globally (Parmar et al., 2010). Stakeholders, defined as those who influence or are influenced by a firm’s activities (Freeman, 1984) or those with legitimate interests in corporate activity (Donaldson and Preston, 1995b), play a pivotal role in this framework. In light of the breadth of actors encompassed by these definitions, considerable effort in stakeholder management research has sought to classify stakeholders to understand how they influence a focal firm (Mitchell et al., 1997). Mitchell et al. (1997) identify three attributes influencing stakeholder salience: (1) Power – the ability to influence outcomes, (2) Legitimacy – alignment with norms and values, and (3) Urgency – the criticality and time sensitivity of stakeholder claims. These attributes shape the strategic attention received by stakeholders.
@@
Building on stakeholder classification, research identifies four broad approaches that firms can adopt: proaction, accommodation, defense, and reaction (Clarkson, 1995; Wartick and Cochran, 1985). Stakeholder influence strategies include coercion and compromise (Frooman, 1999; Frooman and Murrell, 2005). Most research using stakeholder management differentiates between two fundamental motives for firms, including stakeholders, in their decisions, while recognizing that both are always at play to some degree. The first body of stakeholder thinking, what might be termed the instrumental approach, sees stakeholder management as important because managing stakeholder relationships generates significant expected returns for the organisation. A second body of research – the normative approach – emphasises the intrinsic value of taking stakeholders into account and sees stakeholder management as a moral responsibility of contemporary organisations. Instrumental studies show that effective stakeholder management enhances reputation, trust, operational license, and reciprocity (Berman et al., 1999; Choi and Wang, 2009; Clarkson, 1995). Normative research argues that moral imperatives should guide stakeholder considerations, which are often rooted in social contract theory, property rights, and Kantian ethics (Donaldson and Dunfee, 1999; Evan and Freeman, 1988).

### v2

@@ 3 Supply Chain Management
Similarly, the rapid growth of e-commerce has introduced distinct operational complexities, such as the need for efficient last-mile delivery and inventory management in a digital marketplace (Janjevic and Winkenbach, 2020). In the context of e-commerce, Calzavara et al. (2023) investigated different strategies for e-grocery logistics, comparing the costs and performances of various network designs to determine the most efficient approach in response to fluctuating online demand. With reference to the e-commerce sector, Liu et al. (2022) further emphasised that an optimised network configuration may, in certain scenarios, influence the demand pattern itself, potentially leading to a situation where the initial optimisation no longer yields the best performance, thus highlighting the dynamic interplay between logistics design and consumer behaviour and demand. Another paper (Agac et al., 2023) introduced a novel mathematical model using Mixed-Integer Nonlinear Programming for optimising blood supply chain networks, focusing on location, allocation, and routing. A variant of the Genetic Algorithm is applied, and the model’s effectiveness is demonstrated through a case study in the Eastern Black Sea Region of the Turkish Red Crescent.
@@
Purchasing and supply management can contribute to organisational efficiency, risk management and innovation (Axelsson and Håkansson, 1984). In financial terms, this translates into cost reduction, assets optimisation and revenue growth. Growing emphasis is placed on the contribution to non-financial performance dimensions such as environmental and social sustainability, and governance in terms of the accountability, legitimacy and ethics of the organisation’s activities (Quarshie et al., 2016). The exact definition of purchasing and supply management – or, in short: procurement – has seen a shift in emphasis since the 1960s, when the field became subject of academic education and research (Wynstra et al., 2019). The first textbooks focused on differences with consumer buying behaviour and on the tactical and operational purchasing process (e.g., Webster and Wind, 1972). By the 1990s, definitions started to explicate the variety of buying needs, such as the one by Van Weele (1994) that distinguished between direct (Bill of Material) and indirect spend. In the 2000s, definitions put the spotlight on the identification, development and maintenance of relations with suppliers (e.g., Axelsson et al., 2005; Leenders et al., 2006). Consequently, purchasing and supply management is now seen to govern both tactical and operational, more short-term and internally oriented decisions and actions (purchasing) and strategic, more long term and externally oriented decisions and actions (supply), as depicted in figure 6.
@@
The upper, strategic/tactical part of this process model is commonly referred to as ‘sourcing, and is the focus of this section. The bottom part of the model is referred to as ‘purchase to pay’. This particular process model thus captures parts of the procurement process that are often treated separately, dividing between tactical and operational processes (e.g., Monczka et al., 2002, p. 29) versus strategic processes (e.g., Monczka, 1999).
@@
Specification of need pertains to both quantities and qualities. Especially for direct purchase items, need quantities are primarily derived from final demand forecasts (see §4.6). Specification of needs in terms of qualities is typically a cross-functional process, where the procurement department gathers information and preferences from internal customers and technical specialists. This specification process is one of the typical areas where the mandate of the procurement function is being contested. In those organisations where there is no substantial mandate, procurement’s role in the specification-setting process is usually highly limited, and they are relegated to processing orders based on the specifications provided by internal customers. This is very different in organisations where procurement does have a substantial mandate, for instance because of proven capabilities, achieved successes and general management support. In those situations, procurement is involved early on in the process, and has an explicit role in challenging possibly ’gold-plated’ specifications. Also, current and potential suppliers can provide valuable inputs. Giving suppliers influence over the specifications of their offerings positively affects both the efficiency and effectiveness of new product development projects (Suurmond et al., 2020).
@@
In a growing number of organisations, procurement will have the possibility if not the imperative to step up from just purchasing goods and services to becoming a function of real end-to-end value entrepreneurs. Whether the procurement function will be able to fulfill such an increasingly strategic role in a given organisation, depends on multiple factors including the capabilities and attitudes of procurement personnel and managers, its working relationships wither other functions, but surely also its effective application of digital technologies. While successful blockchain applications have become quite sparse by now, the uptake of Articial Intelligence (AI) applications in procurement is rapidly growing, particularly in areas such as negotiations with suppliers of non-critical items (Van Hoek et al., 2022) and other routine tasks such as tender preparations and (initial) bid evaluations (Guida et al., 2023). As with the implementation of digital tools in the early 2000’s, the rationale for these applications is largely that it will enable the procurement function to increase its efficiency in the operational and tactical processes, so that its critical resources can be dedicated to the more strategic processes. It is likely however, that AI will also become instrumental in these processes in the near future. To study these developments and associated challenges, procurement researchers can draw from a diverse and multidisciplinary set of both established and relatively new theories (Tate et al., 2022).
@@
The origins of academic study on the management of buyer-supplier relationships can be briefly traced to the history of economics and specifically contracting through the theories of agency, finance, property rights and stakeholder theory Jensen and Meckling (1976). Following the industrial revolution as firms moved towards ‘new’ mass production techniques in the late 1920s and early 1930s scholars began to study the idea of ‘what is a firm’ and think about where a firm begins and end.
@@
This work was synthesised by the Industrial Marketing and Purchasing (IMP) Group in the early 1980s and 1990s under the leadership of Håkansson (1982) to develop an empirical analysis of firm networks focussing on the relationship interdependencies and typologies of interactions ranging from short-term episodic to long-term ‘deep’ and embedded interactions (Ellram and Ueltschy Murfield, 2019). Collaborative business relationships are viewed as value adding process within the network of organisations or actors. Building on this foundational work we suggest a working definition of buyer-supplier relationships as follows:
@@
The strategic importance of buyer-supplier relationships has been increasingly studied by scholars since the late 1970s and 1980s with the development of Kraljic’s (1983) seminal portfolio analysis tool, positioning spending based on its relative market availability (essentially a measure of risk exposure) and cost or value to suggest management strategies based on a four-quadrant model. Moreover, development in buyer-supplier relationship thinking closely correlates to changes in the economic business environment and particularly in the paradigm changes in manufacturing which can be described as moving from craft, mass, lean and agile (Lamming, 1993) and laterally towards mass-customisation. Combining the portfolio analysis of Kraljic and his analysis of sourcing risk with the literature on contracts, firm boundaries and competitive advantage, led scholars to observe and develop a variety of buyer-supplier positioning and relationship measure concepts, tools and techniques (Bensaou, 1999; Cousins and Spekman, 2003). The practical implications of these portfolios were recognised by consulting firms and industry alike who used them to optimise spend, analyse risk, and manage their supplier relationships to better effect (Cousins and Spekman, 2003).
During this time the business environment moved from local/regional sourcing to international and then global sourcing. The success of Japan following World War II as a major manufacturer employing innovative and novel manufacturing techniques began to change the way Western manufacturers built, designed and sourced their products – moving from one-stop manufacturers to systems integrators, particularly true of the automotive and aerospace industries (Hines et al., 1999). Techniques such as Just-In-Time (JIT) and Total Quality Management (TQM) involved working and coordinating closely with suppliers (Flynn et al., 1995). Practically, this meant manufacturers across many industries had to think much more strategically about how they would work with suppliers to compete. They restructured their supply bases to create tiered levels of suppliers, creating fewer but higher value suppliers, and started to work with their suppliers on new product development projects.
The next trend coincided with the rise of China (and since other parts of East Asia, India and currently Africa) as a low-cost manufacturing environment which encouraged large manufacturing organisations to outsource manufacturing and source products globally. This shift enabled low-cost production but also came with significant risks. More recently, organisations have slowed, or in some cases even reversed, these decisions as they seek to reshore production and diversify the locations of their supply bases in an effort to reduce a broad range of risks, including geographic concentration, geo-political challenges, and sustainability concerns.
@@
Over the last fifty or so years there has been a substantial discourse on all aspects of buyer-supplier relationships creating a rich and varied body of knowledge. It reasons that most problems and questions focused on buyer-supplier relationships are multi-faceted. For scholars in this area, before embarking on research it is important to consider the unit of analysis and then select an appropriate theoretical lens(es). Essentially the research in this area can be classified into three broad but inter-related theoretical areas, economic, strategic and behavioural. Firstly, the economic lens uses the theory of Coase (1937) and specifically Williamson (1985) which is Transaction Cost Economics (TCE). The unit of analysis here is the transaction at product or service level and focuses primarily on the risks of manufacturing within the boundaries of the firm or outside of it – known as market vs hierarchy. The risks are known as transaction costs and contain a variety of characteristics that can potentially be modelled. That said, transaction costs are notoriously difficult to calculate. The second major theoretical lens developed by Prahalad and Hamel (2009) and extended by Barney (1991) is the Resource Based View of the Firm (RBV) – the focus here is on competitive advantage at the level of the firm, creating and managing value added relationships becomes a strategic competitive advantage to the focal firm (Barney, 2012). The seminal work of Dyer and Singh (1998) in this area builds on the previous ideas of the IMP of networks and dyads and theorises the concept of the ‘relational’ view. This paper has been extended (Dyer et al., 2018) to consider the value creation and capture benefits of building alliances and collaborations, viewing this as a dynamic capability. The third and final theoretical lens is ‘social capital’. Social capital can broadly be defined as the network of relationships among people who live and work in a particular society or organisation. Tools such as Social Network Theory (SNA) facilitate the mapping of these network relationships and to understand the strength of these ties and how information and knowledge flows between them (Granovetter, 1983; Wu and Choi, 2005; Kim et al., 2011). Leveraging these networks within buyer-supplier relationships can add significant value to the buyer and supplier organisations by facilitating innovations, learning, new product development and cost reduction and process improvement programmes (Lawson et al., 2009). These three theoretical lenses are not mutually exclusive however scholars tend to select one meta theory, this can be further supported by a secondary theory approach aimed at explaining some of the more nuanced issues within the research question (McIvor, 2009).
@@
The current political, economic, social and technological business environment is extremely difficult to navigate, fraught with supply chain disruptions, aggressive competition, sustainability pressures and move towards de-globalisation and protectionism. It is further compounded with profound changes in the way new technologies such as Artificial Intelligence (AI), Blockchain and the analysis of Big Data are impacting business and supply chain operations. These technologies are creating greater visibility and traceability allowing organisations to analyse and adjust the way they do business offering positive as well as negative consequences (Gligor et al., 2021). Facilitating relationships through the analysis of big data sets, help to analyse and resolve quality problems and maintain standards and improve innovation and problem solving (Rahman et al., 2023). However, it can also drastically reduce personal interactions leading to a lack of personal trust, socialisation and knowledge exchange (Cousins and Menguc, 2006; Lawson et al., 2009). These relational bonds are essential for efficient and effective problem solving and innovation development and knowledge transfer (Lawson et al., 2008).. The impact of new technology while offering many advantages it can also negatively impact interfirm relationships.
@@
Stakeholder management is an increasingly critical aspect of operations and supply chain management (OCSM). Contemporary economic activity is increasingly complex, entailing coordination and collaboration within and between large networks of globally distributed actors that partner in activities through which a given product or service is brought into the market (Co and Barro, 2009; Silvestre et al., 2023). The efficiency and efficacy of the entire supply chain depend on successful coordination among diverse actors (including customers, retailers, financial institutions, governments, logistics providers, distributors, manufacturers, and primary resource extractors and processors) within complex practices and processes (including forecasting, planning, inventory management, scheduling, delivery, just-in-time sourcing, co-location, collaboration, and information sharing; see also Najjar and Yasin, 2023; Silvestre et al., 2023). In this sense, the global economy is increasingly relational and the success of organizations and economies depends on the development and active management of relationships. These relational interdependencies have been shown to be especially significant in the recent geopolitical and pandemic threats to the global economy and society (Devi et al., 2023; Montoya-Torres et al., 2023).
@@
Stakeholder management research proposes that organisational success depends on considering the interests of multiple stakeholders in strategic decision-making (Donaldson and Preston, 1995; Freeman, 1984; Jones, 1995; Laplume et al., 2008; Parmar et al., 2010). Fundamental to stakeholder thinking is the suggestion that “if we adopt as a unit of analysis the relationships between a business and the groups and individuals who can affect or are affected by it, then we have a better chance to” create and distribute value to multiple actors globally (Parmar et al., 2010). Various stakeholder definitions have been proposed in previous research. Freeman (1984) defines stakeholders as individuals or groups that could influence or be influenced by the activities of the firm while Donaldson and Preston (1995, p. 68) define stakeholders as “persons or groups with legitimate interests in procedural and/or substantive aspects of corporate activity”. In light of the breadth of actors encompassed by these definitions, considerable effort in stakeholder management research has sought to classify stakeholders to understand how they influence a focal firm (Mitchell et al., 1997). Mitchell et al. (1997) identified three attributes of firm-stakeholder dyads that raises their salience in management consideration: (1) Power refers to the ability of a stakeholder to influence, produce or effect behaviour, outcomes, processes, objectives, or direction; (2) Legitimacy relates to a claim’s consistency with expected behaviour, structures, values, beliefs, norms, and rules; (3) Urgency refers to the stakeholder and claims that it is critical to the stakeholder and time-sensitive. Combinations of these relational attributes drive the level of strategic attention warranted within a firm’s consideration at a given point in time.
@@
Building on stakeholder classifications, research suggests several broad approaches to stakeholder management that an organisation can adopt, including proaction, accommodation, defence, and reaction (Clarkson, 1995; Wartick and Cochran, 1985). Research has also classified the influence strategies of stakeholders, with Frooman (1999); Frooman and Murrell (2005) proposing two basic influence strategies – coercion and compromise – adopted by stakeholders to influence a focal firm. Most research using stakeholder management differentiates between two fundamental motives for firms, including stakeholders, in their decisions, while recognizing that both are always at play to some degree. The first body of stakeholder thinking, what might be termed the instrumental approach, sees stakeholder management as important because managing stakeholder relationships generates significant expected returns for the organisation. A second body of research –- the normative approach –- emphasises the intrinsic value of taking stakeholders into account and sees stakeholder management as a moral responsibility of contemporary organisations. Studies on instrumental traditions have provided significant evidence that stakeholder management generates organisational benefits mainly through improved reputation, increased trust from stakeholders, an improved ‘license to operate’, and stakeholders acting reciprocally (Berman et al., 1999; Choi and Wang, 2009; Clarkson, 1995; Hillman and Keim, 2001; Jones, 1995). Normative research in the stakeholder tradition has argued that moral imperatives should consider stakeholders, grounding arguments in social contract theory (Donaldson and Dunfee, 1999), property rights (Donaldson and Preston, 1995), and Kantianism (Evan and Freeman, 1988). To some extent, research drawing on signalling theory has sought to bridge the instrumental and normative perspectives by exploring what attributions, inferences, and conclusions stakeholders attach to firms’ breadth and depth of stakeholder management activities (Aronson et al., 2025; Fu et al., 2022).